Consider the three stocks in the following table. Pt represents the price at the end of period t and Qt is the number of shares outstanding. Stock C splits 3:1 during period 2.

P0 Q0 P1 Q1 P2 Q2
A $       115.00 150.00 $       125.00 150.00 $ 128.00 150.00
B $          66.00 400.00 $         72.00 400.00 $   77.00 400.00
C $          35.00 200.00 $         45.00 200.00 $   19.00 600.00

Calculate the price-weighted index return for the period ending in 2. Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

IV(i) = [=1P(i)] / n

IV1 = (125 + 72 + 45) / 3 = 242/3 = 80.67

IV1 = (128 + 77 + 19) / 3 = 224/3 = 74.67

Return = [IV1 / IV0] - 1 = [74.67 / 80.67] - 1 = 0.9256 - 1 = -0.0744, or -7.44%